a rectangular having area of 260 sq mtr,if we increase the width of rectangular by 2 mtr and decrease length by 5 mtr .it becomes a square which also having the same area as rectangle
Answer by mananth(16946)

let length be x
width be y
(x-5) new length
(y+2) new width
Area = (x-5)(y+2)
Area = xy+2x-5y-10=260
in the square, length = width = x
x^2+2x-5x-10=260
x^2-3x-270=0
x^2-18x+15x-270=0
x(x-18)+15(x-18)=0
(x-18)(x+15)=0
x= 18 the length of square
so rectangle width = 18-5=13mtr
length = 18+2=20 mtr
m.ananth@hotmail.ca